# QRazor: Reliable and Effortless 4-bit LLM Quantization by Significant Data Razoring

### v1

@@ 4.3 Decompression-Free Arithmetic Operation
The top part of Figure  3 illustrates the concept of conventional arithmetic operations for QRazor, which involve decompression of operands before computation. In this approach, weight and activation values are decompressed separately to their base precision using logical shift operations, followed by multiplication with a high-precision of 16$\times$8 multiplier. In contrast, the proposed decompression-free arithmetic operation is shown in the bottom part of Figure  3. Here, the operands are directly multiplied using the 4$\times$4 multiplier. The multiplied results are then manipulated by the flag bits using a single 16-bit barrel shifter. This approach eliminates the need for separate decompression steps, thereby enhancing computational efficiency and throughput. The hardware efficiency is further analyzed in Section 5.3.
@@ 5 Experiments
We evaluate our proposed QRazor method on the LLaMA(Touvron et al., 2023a) and LLaMA2(Touvron et al., 2023b) models. Static quantization utilizing the absolute max scaling is implemented for all weights, activations, and KV cache values, with granularity configured as per-channel for weights, per-tensor for activations, and per-head for KV cache values. The SDR technique is applied for 4-bit compression using various group sizes. Weight compression is performed offline, while activations and KV caches are compressed online. Calibration and validation of zero-shot accuracy and perplexity were performed by randomly selecting 128 samples from Wikitext2 for all tasks, with perplexity evaluated on LAMBADA-OpenAI and accuracy on PIQA, ARC, Hellaswag, and Winogrande using lm-evaluation-harness (Gao et al., 2023). Additionally, an ablation study was conducted to assess the impact of group size on quantization accuracy, focusing on the compressed bits for weights and activations. A sequence length of 2048 and a batch size of 1 were used for all tasks during evaluation, and the models provided by meta-llama were employed for the tests.
@@
Table 1: Zero-shot accuracy of W4A4 LLaMA models on five common sense tasks.
@@
Model #Bits Eff. Bits Method PIQA ARC-e ARC-c HellaSwag Winogrande Avg
LLaMA-1-7B FP16 16 - 77.67 76.73 40.13 73.08 68.78 67.28
W4A4 4.03/4 OS+ 62.73 39.98 30.29 44.39 52.96 46.07
W4A4 4 OmniQuant 66.15 45.20 31.14 56.44 53.43 50.47
W4A4 4 QLLM 68.77 53.16 31.14 57.43 56.67 51.84
W4A4 $g16$ 4.25 QRazor 76.61 73.33 37.79 72.26 66.93 65.38
W4A4 $g32$ 4.125 75.24 69.47 36.12 70.98 66.69 63.70
LLaMA-1-13B FP16 16 - 79.16 79.65 43.48 79.07 72.77 70.83
W4A4 4.03/4 OS+ 63.00 40.32 30.38 53.61 51.54 47.77
W4A4 4 OmniQuant 69.69 47.39 33.10 58.96 55.80 52.99
W4A4 4 QLLM 71.38 47.60 34.30 63.70 59.43 55.28
W4A4 $g16$ 4.25 QRazor 77.53 77.72 40.13 76.12 69.53 68.21
W4A4 $g32$ 4.125 76.17 74.56 37.12 74.45 67.64 65.99
LLaMA-2-7B FP16 16 - 79.13 74.39 45.97 76.21 69.30 69.00
W4A4 4.03/4 OS+ 63.11 39.10 28.84 47.31 51.30 45.93
W4A4 4 OmniQuant 65.94 43.94 30.80 53.53 55.09 49.86
W4A4 4 QLLM 67.68 44.40 30.89 58.45 56.59 51.60
W4A4KV4 4/4/4.125 QuaRot(RTN) 72.09 58.88 35.24 65.40 60.69 58.26
W4A4KV4 QuaRot(GPTQ) 76.77 69.87 40.87 72.16 63.77 65.64
W4A4 $g16$ 4.25 75.84 72.63 42.47 72.96 65.67 65.91
W4A4 $g32$ 4.125 QRazor 73.67 70.70 39.46 71.46 64.09 63.88
W4A4KV4 $g16$ 4.25 73.39 70.88 39.80 70.15 64.01 63.65
W4A4KV4 $g32$ 4.125 73.23 67.54 37.46 67.16 60.46 61.17
LLaMA-2-13B FP16 16 - 80.23 77.82 48.76 79.39 72.30 71.70
W4A4 4.03/4 OS+ 64.47 41.46 32.17 59.30 51.38 49.76
W4A4 4 OmniQuant 69.80 47.22 33.79 59.34 55.49 53.13
W4A4 4 QLLM 70.46 48.48 34.39 62.80 55.41 54.31
W4A4KV4 4/4/4.125 QuaRot(RTN) 77.37 70.83 43.69 73.11 67.32 67.16
W4A4KV4 QuaRot(GPTQ) 78.89 72.98 46.59 76.37 70.24 69.79
W4A4 $g16$ 4.25 77.48 75.09 43.47 76.83 70.09 68.59
W4A4 $g32$ 4.125 QRazor 76.82 73.86 41.80 76.30 68.59 67.47
W4A4KV4 $g16$ 4.25 77.80 75.44 44.15 76.06 67.72 68.23
W4A4KV4 $g32$ 4.125 77.09 74.91 44.28 75.06 66.77 67.62
Table 2: Zero-shot accuracy of Mistral-7B & Gemma2-2B models on five common sense tasks.
Zero-shot Accuracy $\uparrow$
Model #Bits Method PIQA ARC-e ARC-c HellaSwag Winogrande Avg
Mistral-7B FP16 Baseline 80.74 81.58 50.16 61.25 74.27 69.60
W4A8 $g16$ QRazor 79.60 80.35 47.15 60.51 72.38 68.00
W4A8 $g32$ 79.76 79.82 46.82 60.11 71.19 67.54
W4A8KV4 $g16$ 79.65 80.18 46.15 59.95 72.35 67.66
W4A8KV4 $g32$ 77.53 76.09 42.24 59.58 66.38 64.36
W4A4 $g16$ QRazor 77.75 77.98 45.48 57.70 69.77 65.74
W4A4 $g32$ 77.80 75.39 43.47 56.20 68.90 64.35
W4A4KV4 $g16$ 77.78 77.34 44.48 58.09 68.56 65.25
W4A4KV4 $g32$ 77.53 76.09 42.24 56.38 66.38 63.72
Gemma-2-2B FP16 Baseline 60.17 37.89 22.41 35.09 51.07 41.32
W4A8 $g16$ QRazor 59.09 38.95 23.41 34.55 52.80 41.76
W4A8 $g32$ 57.83 35.61 22.07 34.79 52.01 40.46
W4A8KV4 $g16$ 59.47 39.12 23.08 34.56 53.35 41.92
W4A8KV4 $g32$ 58.00 36.32 21.07 34.70 51.22 40.26
W4A4 $g16$ QRazor 57.02 38.77 23.08 32.82 51.78 40.69
W4A4 $g32$ 55.71 32.46 20.06 32.13 51.05 38.28
W4A4KV4 $g16$ 57.18 37.54 19.73 32.80 50.51 39.55
W4A4KV4 $g32$ 55.98 33.16 20.74 32.15 47.91 37.99
5.1 LLM Results
LLAMA and LLAMA2: We conduct a comprehensive experiment to evaluate the performance of both LLaMA and LLaMA2 across various model scales. Table 1 presents the overall accuracy results, comparing our QRazor scheme with previous SOTA quantization methods that provide W4A4 precision. Notably, both our approach and QuaRot (Ashkboos et al., 2024) additionally offer 4-bit KV cache quantization. For our QRazor scheme, group sizes of 16 and 32 are evaluated, where groups share the same flag bits. These group sizes yield the same effective bits per data as conventional group-wise quantization schemes with 128 data per group. As previously mentioned, our scheme enables direct low-precision processing, ensuring that smaller group sizes hardly degrade hardware performance, as no dequantization operations are required across groups.
While QLLM (Liu et al., 2024a) preserves the best accuracies in LLaMA models compared to previous methods, it still suffers from significant accuracy degradation, with an average drop exceeding 10%. Additionally, QLLM does not address the quantization of KV caches into a low-precision format, which represents a significant memory bottleneck during long-context LLM inference. In contrast, our QRazor scheme achieves higher accuracy in LLaMA, showing the lowest accuracy drop from the baseline. For LLaMA2 models, QuaRot is compared, showing reasonable results with a marginal accuracy drop in most tasks using W4A4KV4 precision. When comparing the methods equipped with the nearest-to-round (RTN) rounding, commonly applied for both weight and activation compression, QRazor outperforms QuaRot, exhibiting an average accuracy drop of $<$8% in LLaMA2 models. QuaRot suggests that applying GPTQ (Frantar et al., 2023) to the weights could further reduce the accuracy drop. In this case, results with similar or slightly better accuracy are observed compared to our method with RTN applied. Since our scheme is solely influenced by the internal data distribution and is independent of other external criteria, various optimization techniques such as GPTQ can also be applied to our 4-bit model for further enhancements. We leave these potential improvements for future work.
Mistral-7B and Gemma2-2B: In addition to the LLaMA models, we conducted an accuracy analysis on Mistral-7B and Gemma2-2B, achieving strong results across these models, as shown in Table 2. These findings highlight QRazor’s capability to deliver reliable performance across various models and configurations.
5.2 Ablation Studies
W4A8 evaluation. Although QRazor outperforms various 4-bit quantization approaches in task performance, some accuracy degradation is inevitable due to the aggressive low-precision representation. Therefore, as an extension of our scheme, we increase the number of captured salient bits to 8 bits per activation. This section demonstrates the effectiveness of the wider W4A8 version of QRazor, comparing it to QLLM, a SOTA quantization method that also supports the W4A8 setting.
Table 10 presents the accuracy results under the W4A8 precision setting. In the QRazor scheme, the W4A8 configuration demonstrates significantly higher accuracies across various tasks than W4A4. At W4A8 precision, our QRazor consistently outperforms QLLM in most cases. Additionally, we further quantize KV caches to 4 bits, denoted as W4A8KV4, resulting in minimal accuracy loss. When comparing W4A8 to W4A4 in Table 1, QRazor demonstrates less degradation in performance with reduced activation precision compared to QLLM. Notably, our SDR technique captures salient bits at the bit level, enabling finer data preservation than other methods, which proves especially advantageous in more aggressive compression settings.
Compared to another SOTA method, QServe Lin et al. (2024b), QRazor achieves nearly comparable accuracies. However, QServe employs complex rotation operations for quantizing activations, similar to Quarot, which introduces significant computational overhead. This highlights that QRazor is a more efficient technique than QServe, at the W4A8KV4 configuration.
Table 3: Zero-shot accuracy of W4A8 LLaMA2 models on five common sense tasks.
Zero-shot Accuracy $\uparrow$
Model #Bits Method PIQA ARC-e ARC-c HellaSwag Winogrande Avg
@@ 5.2 Ablation Studies
Effect of the SDR group size. This section examines the impact of different group size settings in our scheme. A key feature of our approach is the use of smaller group sizes compared to other methods, as we can perform low-precision computations without the need for decompressing or dequantizing the data, thereby maintaining hardware performance. However, smaller group sizes do result in an increased size of the cached flag data. Meanwhile, other quantization approaches also require additional data, such as FP32 or FP16 scale factors per group, which must be stored and then dequantized after completing the final MAC operations within a group. This implies that each data point gains an additional 0.25 or 0.125 bits ($group$ $size=128$) beyond its quantized bit-width, a metric commonly referred to as the effective bit-width per data.
Table 4 presents the average accuracy of the LLaMA and LLaMA2 models across various tasks when QRazor is applied with different group sizes in W4A4KV4 configuration. As the group size increases, capturing salient bits becomes more challenging due to the larger number of data points sharing the same razoring point. Consequently, significant accuracy degradation is observed in some tasks when tested with a group size of 128, which shares four flag bits within the group, resulting in an effective bit-width of approximately 4.03 bits per data. Based on the average accuracy drop, a group size of 32 or smaller provides minimal accuracy loss across various models, outperforming other methods. Group sizes of 16 and 32 also yield reasonable effective bit-widths of 4.25 and 4.125, respectively, which are comparable to those of group-based quantization methods with a group size of 128. This explains why group sizes of 16 and 32 are selected as the primary configurations for comparison.
@@
Average Accuracy LLaMA-1-7B 67.28 64.79 64.18 61.08 60.06 53.28
LLaMA-1-13B 70.83 67.96 67.55 64.74 63.23 58.53
LLaMA-2-7B 69.00 64.58 63.65 61.17 56.91 47.36
@@
5.3 Hardware Efficiency
@@
  FP 16$\times$16 MAC   INT 16$\times$8 MAC   INT 8$\times$8 MAC   INT 4$\times$4 Proposed
@@ 6 Conclusion
In this paper, we introduced QRazor, a post-training quantization (PTQ) method that enables reliable 4-bit quantization for large language models (LLMs) through two key stages: quantization and compression. Our approach effectively balances accuracy and efficiency by first quantizing weights, activations, and KV caches to 8-bit and 16-bit integer formats, followed by compressing them using the significant data razoring (SDR) technique, which preserves salient bits for low-bit computations. Experimental results on LLaMA and LLaMA2 models demonstrate QRazor’s reliable performance compared to state-of-the-art methods, particularly in scenarios such as W4A4 and W4A4KV4, with minimal accuracy degradation when using smaller group sizes. Additionally, our hardware-optimized, decompression-free arithmetic unit significantly reduces area and power consumption, making QRazor ideal for deployment in resource-constrained environments. Overall, QRazor presents an efficient and scalable solution for low-bit quantization in LLMs, winning both software and hardware optimizations to enhance system performance.
@@ Appendix A Appendix
A.1 Precision Analysis of the Quantization Stage
As discussed in Section 4.1, outliers pose a significant challenge in low-bit quantization of LLMs, particularly for activations, which exhibit a much broader dynamic range compared to weights or KV caches. Our observations, consistent with previous research (van Baalen et al., 2023), indicate that activations require quantization to higher precision, such as 16-bit integers, to capture this wide range without sacrificing accuracy. On the other hand, weights and KV caches can be effectively represented using 8-bit integers without significantly affecting LLM accuracy. This finding underscores the necessity of selecting appropriate bit-widths during the quantization stage to effectively capture the full range of activation outliers, ensuring accurate representation without compromising model performance.
To validate these requirements, we evaluated LLM accuracy across three quantization scenarios: W8A8 (8-bit weights and 8-bit activations), W8A16, and W8A16KV8. Our results confirm that the W8A16 and W8A16KV8 scenarios achieve nearly identical accuracies to their FP16 counterparts across various LLaMA tasks, while W8A8 results in significant accuracy drops. These findings demonstrate that 16-bit integers are crucial for capturing the outlier-prone dynamic range of activations, whereas 8-bit integers suffice for weights and KV caches. Note that our razoring scheme involves a delicate process of detecting salient bits at the bit level, but it must be applied to the base integer data with minimal information loss. Selecting the appropriate base precision during the quantization stage is, therefore, crucial to maintaining data accuracy.
Table 6: Zero-shot accuracy of different base precision settings.
Zero-shot Accuracy $\uparrow$
Model #Bits Method PIQA ARC-e ARC-c HellaSwag Winogrande
LLaMA-2-7B FP16 - 79.13 74.39 45.97 76.21 69.30
W8A8 71.89 65.99 33.11 65.11 64.33
W8A16 Absmax Quantization 77.69 74.05 42.43 75.97 69.30
W8A16KV8 78.07 74.20 42.41 76.26 69.30
LLaMA-2-13B FP16 - 80.23 77.82 48.76 79.39 72.30
W8A8 68.99 40.35 27.09 57.26 50.99
W8A16 Absmax Quantization 78.94 76.47 45.14 78.92 71.27
W8A16KV8 78.73 76.38 44.81 79.03 71.51
A.2 Importance of Weight Quantization

### v2

@@ 4.3 Decompression-Free Arithmetic Operation
The top part of Figure 3 illustrates the concept of conventional arithmetic operations for QRazor, which involve decompression of operands before computation. In this approach, weight and activation values are decompressed separately to their base precision using logical shift operations, followed by multiplication with a high-precision of 16$\times$8 multiplier. In contrast, the proposed decompression-free arithmetic operation is shown in the bottom part of Figure 3. Here, the operands are directly multiplied using the 4$\times$4 multiplier. The multiplied results are then manipulated by the flag bits using a single 16-bit barrel shifter. This approach eliminates the need for separate decompression steps, thereby enhancing computational efficiency and throughput. The hardware efficiency is further analyzed in Section 5.4.
@@ 5.1 Setup
We evaluate our proposed QRazor method on the LLaMA2(Touvron et al., 2023) and LLaMA3(Grattafiori et al., 2024) models. Static quantization utilizing the absolute max scaling is implemented for all weights, activations, and KV cache values, with granularity configured as per-channel for weights while per-tensor for activations and KV cache values. The SDR technique is applied for 4-bit compression using various group sizes. Weight compression is performed offline, while activations and KV caches are compressed online. Calibration and validation of zero-shot accuracy and perplexity were performed by randomly selecting 128 samples from Wikitext2 for all tasks, with perplexity evaluated on the same task while accuracy on PIQA, ARC, Hellaswag, and Winogrande using lm-evaluation-harness (Gao et al., 2023). Additionally, an ablation study was conducted to assess the impact of group size on quantization accuracy, focusing on the compressed bits for weights and activations. A sequence length of 2048 and a batch size of 1 were used for all tasks during evaluation, and the models provided by meta-llama and mistralai-mistral were employed for the tests.
5.2 LLM Results
LLaMA-2 and LLaMA-3: We conduct comprehensive experiments to evaluate the performance of both LLaMA-2 and LLaMA-3 across various model scales. Table 2 presents the overall accuracy results, comparing our QRazor scheme with previous SOTA quantization methods that provide W4A4 precision. Notably, both our approach and QuaRot (Ashkboos et al., 2024) additionally offer 4-bit KV cache quantization. For our QRazor scheme, group sizes of 16 and 32 are evaluated, where groups share the same flag bits. These group sizes yield the same effective bits per data as conventional group-wise quantization schemes with 128 data per group. As previously mentioned, our scheme enables direct low-precision processing, ensuring that smaller group sizes hardly degrade hardware performance, as no dequantization operations are required across groups.
@@ 5.2 LLM Results
Table 2: Zero-shot accuracy of LLaMA families and Mistral-7B on five common sense tasks.
Perplexity $\downarrow$ Zero-shot Accuracy $\uparrow$
Model #Bits Eff. Bits Method Wikitext2 PIQA ARC-e ARC-c HellaSwag Winogrande Avg
LLaMA-2-7B FP16 16 - 5.47 79.13 74.39 45.97 76.21 69.30 69.00
W4A4 4.03/4 OS+ - 63.11 39.10 28.84 47.31 51.30 45.93
W4A4 4 OmniQuant 14.61 65.94 43.94 30.80 53.53 55.09 49.86
W4A4 4 QLLM 11.75 67.68 44.40 30.89 58.45 56.59 51.60
W4A4KV4 4/4/4.125 QuaRot(RTN) 8.37 72.09 58.88 35.24 65.40 60.69 58.26
W4A4KV4 QuaRot(GPTQ) 6.10 76.77 69.87 40.87 72.16 63.77 65.64
W4A4 $g16$ 4.25 6.45 75.84 72.63 42.47 72.96 65.67 65.91
W4A4 $g32$ 4.125 QRazor 7.03 73.67 70.70 39.46 71.46 64.09 63.88
W4A4KV4 $g16$ 4.25 7.62 73.39 70.88 39.80 70.15 64.01 63.65
W4A4KV4 $g32$ 4.125 8.14 73.23 67.54 37.46 67.16 60.46 61.17
LLaMA-2-13B FP16 16 - 4.88 80.23 77.82 48.76 79.39 72.30 71.70
W4A4 4.03/4 OS+ - 64.47 41.46 32.17 59.30 51.38 49.76
W4A4 4 OmniQuant 12.28 69.80 47.22 33.79 59.34 55.49 53.13
W4A4 4 QLLM 9.09 70.46 48.48 34.39 62.80 55.41 54.31
W4A4KV4 4/4/4.125 QuaRot(RTN) 6.09 77.37 70.83 43.69 73.11 67.32 67.16
W4A4KV4 QuaRot(GPTQ) 5.40 78.89 72.98 46.59 76.37 70.24 69.79
W4A4 $g16$ 4.25 5.37 77.48 75.09 43.47 76.83 70.09 68.59
W4A4 $g32$ 4.125 QRazor 5.54 76.82 73.86 41.80 76.30 68.59 67.47
W4A4KV4 $g16$ 4.25 5.49 77.80 75.44 44.15 76.06 67.72 68.23
W4A4KV4 $g32$ 4.125 5.79 77.09 74.91 44.28 75.06 66.77 67.62
LLaMA-3-8B FP16 16 - 6.14 80.65 80.18 50.84 79.18s 73.01 72.77
W4A4KV4 4/4/4.125 QuaRot(RTN) - - - - - - -
W4A4KV4 QuaRot(GPTQ) 8.16 75.14 68.01 43.34 72.94 72.77 65.18
W4A4 $g16$ 4.25 6.49 78.42 79.65 46.82 77.28 71.90 70.81
W4A4 $g32$ 4.125 QRazor 7.05 78.24 77.89 45.82 76.93 71.40 70.06
W4A4KV4 $g16$ 4.25 7.57 77.42 77.19 44.15 75.81 69.30 68.77
W4A4KV4 $g32$ 4.125 8.26 76.77 75.51 42.81 74.05 68.75 67.58
Mistral-7B FP16 16 - 5.42 80.74 81.58 50.16 61.25 74.27 69.60
W4A4 $g16$ 4.25 6.06 77.75 77.98 45.48 57.70 69.77 65.74
W4A4 $g32$ 4.125 QRazor 6.24 77.80 75.39 43.47 56.20 68.90 64.35
W4A4KV4 $g16$ 4.25 6.18 77.78 77.34 44.48 58.09 68.56 65.25
W4A4KV4 $g32$ 4.125 6.93 77.53 76.09 42.24 56.38 66.38 63.72
While QLLM (Liu et al., 2024a) preserves the acknowledgeable accuracies in LLaMA-2 models compared to previous methods, it still suffers from significant accuracy degradation. Additionally, QLLM does not address the quantization of KV caches into a low-precision format, which represents a significant memory bottleneck during long-context LLM inference. In contrast, our QRazor compresses KV caches to 4-bit while achieving higher accuracy than QLLM in LLaMA-2 models. For LLaMA-2 models, QuaRot is also compared, showing reasonable results with a marginal accuracy drop in most tasks using W4A4KV4 precision. When comparing the methods equipped with the rounding-to-nearest (RTN), commonly applied for both weight and activation compression, QRazor outperforms QuaRot, exhibiting an average accuracy drop of $<$8% in LLaMA-2 models. QuaRot suggests that applying GPTQ (Frantar et al., 2023) to the weights could further reduce the accuracy drop. In this case, results with similar or slightly better accuracy are observed compared to our QRazor with RTN applied. However, for the LLaMA-3-8B model, QRazor provides better performances than QuaRot regardless of their weight quantization approaches. Since our scheme is solely influenced by the internal data distribution and is independent of other external criteria, various optimization techniques such as GPTQ can also be applied to our 4-bit model for further enhancements. We leave these potential improvements for future work.
Mistral-7B: In addition to the LLaMA models, we conducted an accuracy analysis on Mistral-7B achieving strong results across these models, as shown in Table 2. These findings highlight QRazor’s capability to deliver reliable performance across various models and configurations.
5.3 Ablation Studies
W4A8 evaluation. Although QRazor outperforms various 4-bit quantization approaches in task performance, some accuracy degradation is inevitable due to the aggressive low-precision representation. Therefore, we evalutate W4A8 version of QRazor by increasing the number of captured salient bits to 8 per activation.
Table 3: Zero-shot accuracies of W4A8 LLaMA2 models.
@@ 5.3 Ablation Studies
Model #Bits Method PQ AE AC HS WG Avg
@@
Table 3 presents the accuracy results under the W4A8 precision setting. The W4A8 configuration demonstrates significantly higher accuracies across various tasks than W4A4. At W4A8 precision, our QRazor consistently outperforms QLLM in most cases. Additionally, we further quantize KV caches to 4 bits, denoted as W4A8KV4, resulting in minimal accuracy loss. When comparing W4A8 to W4A4 in Table 2, QRazor demonstrates less degradation in performance with reduced activation precision compared to QLLM. Notably, our SDR technique captures salient bits at the bit level, enabling finer data preservation than other methods, which proves especially advantageous in more aggressive compression settings.
Compared to another SOTA method, QServe (Lin et al., 2024b), QRazor achieves nearly comparable accuracies. However, QServe employs complex rotation operations for quantizing activations, similar to Quarot, which introduces significant computational overhead. This highlights that QRazor is a more efficient technique than QServe, at the W4A8KV4 configuration.
@@
Avg. Accuracy LLaMA-2-7B 69.00 64.58 63.65 61.17 56.91 47.36
@@
LLaMA-3-8B 72.57 70.43 68.77 67.58 65.79 60.44
Effect of the SDR group size. A key feature of our approach is the use of smaller group sizes compared to other methods, as we can perform low-precision computations without the need for dequantization, thereby maintaining hardware performance. However, smaller group sizes do result in an increased size of the cached flag data. Meanwhile, other group-based quantization approaches require additional data due to the scale factors generated for each group. For example, FP32 and FP16 scale factors add an overhead of 0.25 and 0.125 bits per value, respectively, when the group size is 128. We evaluate QRazor under various group size settings to assess its efficiency.
Table 4 presents the average accuracy of LLaMA models across various tasks when QRazor is applied with different group sizes in W4A4KV4 configuration. As the group size increases, capturing salient bits becomes more challenging due to the larger number of data points sharing the same razoring point. Consequently, significant accuracy degradation is observed in some tasks when tested with a group size of 128, which shares four flag bits within the group, resulting in an effective bit-width of approximately 4.03 bits per data. Based on the average accuracy drop, a group size of 32 or smaller provides minimal accuracy loss across various models, outperforming other methods. Group sizes of 16 and 32 also yield reasonable effective bit-widths of 4.25 and 4.125, respectively, which are comparable to those of group-based quantization methods with a group size of 128. This explains why group sizes of 16 and 32 are selected as the primary configurations for comparison.
@@
   FP 16$\times$16    INT 16$\times$8    INT 8$\times$8    INT 4$\times$4
MAC MAC MAC Proposed
@@ 6 Conclusion
We introduced QRazor, a post-training quantization (PTQ) method that enables reliable 4-bit quantization for large language models (LLMs) through two key stages: quantization and compression. Our approach effectively balances accuracy and efficiency by first quantizing weights, activations, and KV caches to 8-bit and 16-bit integer formats, followed by compressing them using the significant data razoring (SDR) technique, which preserves salient bits for low-bit computations. Experimental results on various LLMs demonstrate QRazor’s reliable performance compared to state-of-the-art methods in scenarios such as W4A4 and W4A4KV4. Additionally, our hardware-optimized, decompression-free arithmetic unit significantly reduces area and power consumption, making QRazor ideal for deployment in resource-constrained environments.
@@ Appendix A Appendix
A.1 Importance of the Weight Quantization